OpenStudy (anonymous):

The intersection of the plane z=3 with elliptical cylinder x^2/9+y^2/16=1, find the appropriate parametrization, can you help I hit a wall with this section...lost...yay me

OpenStudy (anonymous):

On the assumption that 'appropriate parametrization' means you are wanting a single parameter to map the locus of a point along the curve ie.\[(x, y,3)\]where x and y are related by :\[\frac{x^2}{9} + \frac{y^2}{16} = 1\]then try :\[x=3cos(t)\]\[y=4sin(t)\]\[z=3\]where\[t\in \Real\]

OpenStudy (anonymous):

omg is that all they're askin for? This section...gonna kill me! thank!